OpenStudy (southernbell82):

Urgent Please Help!!! Last Project of the year and school ends today and I will fail Geometry . Can anyone help??

OpenStudy (southernbell82):

I have completed some of the project but, I am completely lost on the rest of it!

OpenStudy (smartnerd1111):

hi

OpenStudy (amtran_bus):

We need to know what the project is. :)

OpenStudy (smartnerd1111):

yeah what is the projedct about

OpenStudy (southernbell82):

Hey! I am attaching it!

OpenStudy (southernbell82):

I am going to fail if this is not done by 11:59 :(

OpenStudy (southernbell82):

does anyone know how to do the rest of this??

OpenStudy (southernbell82):

@AmTran_Bus @agent0smith @Vocaloid

Vocaloid (vocaloid):

ok so i'm at the packaging shapes part

Vocaloid (vocaloid):

for cylinders and spheres, an example of a pro is that the volume is high compared to surface area

Vocaloid (vocaloid):

cones and pyramids...

Vocaloid (vocaloid):

hm.

Vocaloid (vocaloid):

any ideas for pros?

OpenStudy (southernbell82):

I completely lost and I am going to fail!

OpenStudy (amtran_bus):

@Vocaloid is on fire!

OpenStudy (amtran_bus):

Why did you wait till last minute?????????

Vocaloid (vocaloid):

can you think of any reasons why a pyramid or cone shape might be a good package idea?

OpenStudy (southernbell82):

I was working on other projects :(

OpenStudy (southernbell82):

No. Only that they would be easier to pour from

Vocaloid (vocaloid):

that's good!

Vocaloid (vocaloid):

you can write that

Vocaloid (vocaloid):

for cons, you could say that they can't be stacked

Vocaloid (vocaloid):

well not easily

OpenStudy (wolf1728):

Cones, Pyramids AND spheres can't be tightly packed

OpenStudy (southernbell82):

Okay thanks how about pros for those??

OpenStudy (southernbell82):

Cylinder and Sphere?

OpenStudy (wolf1728):

Sphere would be cheapest in terms of surface area to volume

Vocaloid (vocaloid):

well, pyramids and cones you mentioned before (pouring?) spheres have high volume/surface area

Vocaloid (vocaloid):

I just went over cylinder at the beginning

OpenStudy (southernbell82):

So Sphere would be cheapest to make?

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

now all that's left is triangular prisms and rectangular prisms

OpenStudy (southernbell82):

Yep for that part!

OpenStudy (southernbell82):

What about cube pro/con?

Vocaloid (vocaloid):

oh for cubes well

OpenStudy (wolf1728):

Cube con - inconvenient shape - doesn't stack well in store or at home - hard to hold with one hand.

Vocaloid (vocaloid):

pros for cube is high volume/surface area ratio

OpenStudy (southernbell82):

Thank Y'all How about Triangle Prism Pro/Con?

OpenStudy (wolf1728):

Triangle prism pro - doesn't fall over (if manufactured with a low height)

Vocaloid (vocaloid):

oh

Vocaloid (vocaloid):

oh right, triangles are actually more stable

OpenStudy (southernbell82):

Cool Thanks!! Help me pick one that would cost the least to make!

OpenStudy (wolf1728):

Look at this

OpenStudy (wolf1728):

In terms of cardboard cost sphere is cheapest

OpenStudy (southernbell82):

What measurements should I use on the sphere to equal the first part of the file I attached?

OpenStudy (southernbell82):

I am on the one right before the last question. Right above the net.

Vocaloid (vocaloid):

your volume was 240 inch^2

OpenStudy (southernbell82):

This shape must hold same amount of cereal but reduce costs.

OpenStudy (southernbell82):

Okay!

Vocaloid (vocaloid):

you solve 240 = a^3 for a

Vocaloid (vocaloid):

any ideas how?

OpenStudy (wolf1728):

sphere volume = (4/3) PI* r^3

Vocaloid (vocaloid):

her work sheet says she chose a cube though?

Vocaloid (vocaloid):

unless you wanna switch, that's ok too

OpenStudy (wolf1728):

Okay - sorry :-(

OpenStudy (southernbell82):

Will the cube be cheap also?

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

the formula will be easier to calculate

Vocaloid (vocaloid):

also the cube won't roll away randomly

OpenStudy (southernbell82):

So what do I put in the second section that says Surface area? And Volume?

Vocaloid (vocaloid):

volume first

Vocaloid (vocaloid):

if 240 = a^3, what does a = ?

OpenStudy (southernbell82):

6.12?

Vocaloid (vocaloid):

6.21

OpenStudy (southernbell82):

That's what I meant to type.

Vocaloid (vocaloid):

the dimensions of a cube are equal so

OpenStudy (wolf1728):

6.21 is it

Vocaloid (vocaloid):

*6.21

Vocaloid (vocaloid):

for all

OpenStudy (southernbell82):

So I put that in the surface area box?

Vocaloid (vocaloid):

volume

Vocaloid (vocaloid):

not surface area

OpenStudy (wolf1728):

Surface area = 6*21*6.21*6

Vocaloid (vocaloid):

right

Vocaloid (vocaloid):

just to clarify

Vocaloid (vocaloid):

volume = 6.21^3 = 240 surface area = 6.21 * 6.21 * 6 = 231

Vocaloid (vocaloid):

now for the part where it says "Prove that your new design holds the same amount but can be manufactured more cheaply."

Vocaloid (vocaloid):

just show that the new surface area (231 inch^2 is less than the old surface area 292) but the volumes are both equal (240)

Vocaloid (vocaloid):

ready to keep going?

OpenStudy (southernbell82):

Yes I am!! You are a lifesaver!!! :)

Vocaloid (vocaloid):

the part where it says draw a net

OpenStudy (wolf1728):

Yes vocaloid has done quite well!!

OpenStudy (southernbell82):

What numbers do I put on it?

Vocaloid (vocaloid):

since all the sides are the same, just label each side 6.21

OpenStudy (southernbell82):

Oh My Goodness!! I will pass Geometry because of You!! :)

Vocaloid (vocaloid):